True or False.

If "f(x) = 1\/\u221ax" , then "f'(0)" does not exist.


1
Expert's answer
2021-05-20T11:34:24-0400

"f(x)=\dfrac {1}{\sqrt x}" , now to check that derivative exist or not , we know that to check the derivative of function exist or not , we apply formula -


"f^{'}(x)= _{{lim\ h}{\to}0}\dfrac{f(x+h)-f(x)}{h}"

, now "_{limh\to0}" "\dfrac {\dfrac {1}{\sqrt {x+h}}-\dfrac{1}{\sqrt x}}{h}"


"=_{limh\to0}" "\dfrac {{ \dfrac {{\sqrt x}-{\sqrt{ x+h}}}{{\sqrt x}{\sqrt {x+h}}}}}{h}"


"=_{limh\to0}" "\dfrac{x+{\sqrt x}{\sqrt {x+h}}-{\sqrt{x+h}{\sqrt {x}}-(x+h)}}{h{\sqrt{x}{\sqrt {x+h}}}({\sqrt{x}+\sqrt{x+h)}}}" "=\dfrac {1}{x^{\dfrac {5}{2}}}"


hence we can clearly see that f'(0), of a function tends to infinity hence f'(0) does not exist .